You invested $27,000 into two accounts. One paying 3% annual interest, the other paying 4% annual interest. If the total interest earned for the year was $870, how much was invested into each account at each rate?
Explain, and please show work?

Answer by richwmiller(17219)

x+y=27000,
0.03*x+0.04*y=870
x=27000-y
0.03*(27000-y)+0.04*y=870
810-0.03y+0.04*y=870
0.01*y=60
y= 6000
x=27000-y
x=21000
check
0.03*21000+0.04*6000=870
630+240=870
870=870
ok